OpenStudy (anonymous):

how to integrate 2t[(2+4t^2)^1/2] dt?

OpenStudy (amorfide):

you will have to use integration by parts I believe

OpenStudy (irishboy123):

nah. **differentiate** (2+4t^2)^3/2 and see what you get. work from there.

OpenStudy (amorfide):

\[\int\limits_{a}^{b}u(t)v'(t)=u(t)v(t) - \int\limits_{a}^{b}u'(t)v(t)\]

OpenStudy (anonymous):

Thanks amorfide. I'll try that.

OpenStudy (anonymous):

is there anyway to check if my answer was right?

OpenStudy (welshfella):

yes differentiate your answer to see if it comes back to the function in the question

OpenStudy (anonymous):

how do u choose which one is u and v?

OpenStudy (welshfella):

thats a good question - you want to make the integration u'(t)v(t) relatively easy

OpenStudy (welshfella):

http://www.wolframalpha.com/input/?i=integrate+2t%5B+%282+%2B+4t%5E2%29%5E%281%2F2%29%5D+dt wolframalpha doesn't like it!

OpenStudy (welshfella):

maybe you have the question wrong?

OpenStudy (phi):

I would let \[ u = 4t^2+2 \\ du = 8t~ dt \]

OpenStudy (welshfella):

yes that substitution seems promising

OpenStudy (phi):

thus \[ dt = \frac{du}{8t} \] use that and u in your original equation

OpenStudy (welshfella):

you would have to convert 2t to an expression in u

OpenStudy (welshfella):

- no maybe not

OpenStudy (welshfella):

the 2t would cancel out

OpenStudy (anonymous):

i found it! Thanks so much phi! Will give u a medal xD

OpenStudy (welshfella):